OpenStudy (anonymous):

tanø * cotø / cscø = sinø Need on verifying the identity please... Step by step..?

OpenStudy (whpalmer4):

Break it down into the basic trig functions: \[\tan \theta = \frac{\sin\theta}{\cos\theta}\]\[\cot \theta = \frac{\cos\theta}{\sin\theta}\]\[\csc \theta = \frac{1}{\sin\theta}\]

OpenStudy (whpalmer4):

\[\tan\theta*\frac{\cot\theta}{\csc\theta} = \sin\theta\]\[\frac{\sin\theta}{\cos\theta}*\frac{\frac{\cos\theta}{\sin\theta}}{\frac{1}{\sin\theta}}=\sin\theta\]Do the cancellation and what do you get?